OpenStudy (anonymous):

can you plz show work? needs to be factored 18x^2+48x=-32

OpenStudy (anonymous):

\[18x^2 + 48x = -32\] \[18x^2 + 48x+32=0\]

OpenStudy (anonymous):

(6x + 8)(3x + 4)

OpenStudy (anonymous):

Sorry (6x + 8)(3x + 4) = 0

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

\[2(9x^2+24x+16)\] \[2(3x+4)(3x+4)\] \[2(3x+4)^2\]

OpenStudy (anonymous):

=0